Is fringe width uniform in YDSE ?
Plane screen.



which width will be grater between 10th and 11th or..that of between 1st and 2nd ? give me a equation describing it

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
\
Yes in YDSE all periphery widths are equivalent. 
For nth and mth splendid edges, 
∆x=(n-m+1/2) where =D/d where D=Distance amongst opening and screen, d = separate between the openings. 
Between mth brilliant and nth dim periphery 
∆x=(m-n-1/2)  
 is sure. The tenth and eleventh periphery width will be equivalent to that of first and second in light of the fact that the condition to be connected for them is same i.e the last from the over 2 conditions.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)